OpenStudy (anonymous):

Find the prime –power representation of each of the following numbers: 84

hartnn (hartnn):

start dividing your number with smallest prime number, (which is...?)

OpenStudy (anonymous):

1

hartnn (hartnn):

1 is neither prime nor composite.

OpenStudy (anonymous):

oh.. then 3

hartnn (hartnn):

where id 2 go ?? :O :P

hartnn (hartnn):

*did

hartnn (hartnn):

its basics to know that 2 is smallest prime.... start dividing your number with 2, what u get ?

OpenStudy (anonymous):

oh, thats right..... 2

OpenStudy (anonymous):

42

hartnn (hartnn):

yes, 84 =2*42 is 42 still divisible by 2 ? if yes, then divide again

hartnn (hartnn):

continue dividing until the numbe you get is no longer divisible by 2

OpenStudy (anonymous):

21

hartnn (hartnn):

so, 84 = 2*2*21 = 2^2 *21 now since, 21 is not divisible by 2, we go for next smallest prime after 2 (which is...?)

OpenStudy (anonymous):

3

hartnn (hartnn):

divide 21 by 3 then

OpenStudy (anonymous):

7

hartnn (hartnn):

so, 84 = 2^2 * 3*7 right ?

OpenStudy (anonymous):

yes

hartnn (hartnn):

since these are prime numbers and cannot be divided further, your prime –power representation will be \(84 =2^2 3^17^1\)

OpenStudy (skullpatrol):

Prime factors of positive integers are found by using the primes in order as divisors. The prime factorization of a positive integer is the expression of the integer as a product of prime factors.

OpenStudy (anonymous):

ok. So the number 432 would be 2^4 3^3

hartnn (hartnn):

yes, thats correct! good work :)

OpenStudy (anonymous):

Yay! :)

OpenStudy (anonymous):

what about the number 109?

hartnn (hartnn):

109 is itself a prime number

OpenStudy (anonymous):

oh, so what would I put? just leave it 109?

hartnn (hartnn):

yeah, just 109 or 109^1

OpenStudy (anonymous):

ok. What about 17,017?